### v1

# Inferring Belief States in Partially-Observable Human-Robot Teams

Abstract
We investigate the real-time estimation of human situation awareness using observations from a robot teammate with limited visibility. In human factors and human-autonomy teaming, it is recognized that individuals navigate their environments using an internal mental simulation, or mental model. The mental model informs cognitive processes including situation awareness, contextual reasoning, and task planning. In teaming domains, the mental model includes a team model of each teammate’s beliefs and capabilities, enabling fluent teamwork without the need for explicit communication. However, little work has applied team models to human-robot teaming. We compare the performance of two current methods at estimating user situation awareness over varying visibility conditions. Our results indicate that the methods are largely resilient to low-visibility conditions in our domain, however opportunities exist to improve their overall performance.
I Introduction
The human factors community recognizes that when individuals conduct taskwork, they construct an internal simulation termed a mental model. The mental model broadly informs the individual’s abstract understanding of the environment, such as contextual information, the status of the task at hand, contingencies, and actionable plans [3, 17].
In collaborative tasks, the mental model expands to include a team model incorporating a prediction teammate beliefs, characteristics, capabilities, and roles [23, 17]. By constructing a team model in addition to their own mental model, individuals apply the information they believe their teammates know – and do not know – in their own plans. This allows individuals to work proactively by anticipating teammate needs without explicit communication [22, 14].
Recent research interest has sought to apply team models for human-robot and human-autonomy teaming [11] [15]. Researchers in the human-robot interaction (HRI) community anticipate that an inferred understanding of human teammate beliefs can support a variety of robot planning, comprehension, and decision-making goals [21]. Related work has proposed frameworks for constructing team models [19] [2], however few work have evaluated these frameworks in human-robot teaming domains.
Human factors researchers often use situation awareness as a proxy for pertinent components of the mental model [8, 21]. Situation awareness is an individual’s understanding of the current environment along three levels: the raw world state (e.g., object locations), contextual information (e.g., inferred properties and relationships between objects), and future projection (e.g., plans and dynamic models). In practice, situation awareness is typically measured in-situ by periodically halting the environment and asking the user questions about the task’s current state [9].
Our objective is to evaluate two models that enable robots to predict human teammates’ belief states. We define the belief state as measurable information of the teammate’s situation awareness. We split the problem into two components – estimating the user’s awareness of world state elements, and estimating how that information is used to answer Level 1 (world state) and Level 2 (contextual) situation awareness questions. Each model addresses the second component by reasoning over a scene graph of the user’s estimated knowledge of the current task.
Concretely, this paper contributes the following:
  1. 1.
    An online human-robot teaming domain and situation awareness dataset featuring partial observability and two belief state prediction models – a logical predicates model and a large language model (LLM).
  2. 2.
    Findings that the belief state prediction models were resilient to low-observability conditions and moderately agreed with human user responses.
Our results indicate that both models were successful at inferring user situation awareness. Moreover, both models maintained their performance under varying visibility conditions. The LLM performed on par with the hand-crafted logical predicates model, indicating that the LLM can be leveraged to create rules for logical predicates models, or encapsulate the process entirely.
All data and source code for this paper is available at https://jackkolb.com/tmm-hri.
II Related Works
The HRI community has primarily utilized mental models with task-dependent factors aligned to the three levels of situation awareness. For example, in a process-driven task (e.g., manufacturing, housework, spaceflight), the mental model may contain a representation of the overall procedure (Level 1), where in the procedure the environment is currently (Level 2), and what that implies for the immediate future (Level 3). Contextual information is derived from the world state, which supports higher-level reasoning about the future of the environment and task. In addition, mental models can draw upon prior experiences to form a stronger model of the risks, failure modes, and practical effects of different actions and task strategies [22].
Mental models are also believed to play an important role in teaming [22]. In human-human teams, people predict the mental model of their human teammates, the team model. If an individual believes their teammate’s mental model is incorrect, or suspects their own mental model is outdated, they adapt their plans or communicate with their teammate accordingly. Related work has found that a strong team model is often important for high-performing teams [14].
A team model also facilitates theory of mind [5]. Theory of mind is an agent’s understanding of another agent’s beliefs, such as the other’s knowledge, intentions, emotions, and opinions. Theory of mind is generally studied in the context of education, where instructors aim to shape the belief states of their students, and lends itself to orders of intentionality (“I believe that you know X”). Theory of mind is inherently team-centric, and has been studied in several close-proximity human-robot collaboration domains [6, 20].
Most related work on mental models has focused on humans and human-human teams, however in recent years there has been interest in expanding mental model theory to human-robot teams [17, 12, 10, 13]. Work in this space has been challenged by the extraordinary complexity of representing human cognition, which remains an existential cognitive science problem. This has led researchers to use simplified domains and team models. This work uses a domain that more closely represents real-world human-robot teaming applications.
II-A Logical Predicates Representation
The most prevalent representation of mental models in the HRI literature uses logical predicates, or sets of known logical truths [7, 18, 10, 19]. Each predicate is an observed feature of the environment state taken as an absolute truth, e.g., the location of an object and its properties. Complex functions are hand-crafted to derive contextual information or model how the environment will change over time. When predicting the mental model of a teammate, a robot agent can represent the teammate’s mental model using the same logical predicates framework, and filter the observational inputs by the workspace region visible to the teammate.
The logical predicates model is attractive for environments with low observational uncertainty. When there is little room for the human teammate to misunderstand their observations, and the robot has full observability, the teammate’s mental model is a subset of the robot’s mental model. With these assumptions, a rigorously-designed logical predicates model can capture many if not all the important aspects of the environment and answer a wide range of inference questions.
However, the robot’s mental model can be misled when it does not contain all information available to the teammate’s mental model, for example, in partially observable conditions. As the logical predicates model is fact-based and cannot model uncertainty, hard-coding reasoning rules to resolve conflicting information can negatively impact downstream planning and reasoning.
Research opportunities currently exist to consider uncertainty within mental models. Specifically, the literature identifies three areas for error: misperception of events, miscomprehension of perceived events, and imperfect mapping of perceived events to the updated mental model [2].
In this work we assess whether an LLM can match a logical predicates model at inference over a scene graph, challenging the need for hand-crafted inference rules.
II-B Markov Decision Process Representation
Nikolaidis et al. proposed representing the team model as a partially-observable Markov decision process (POMDP), where the environment is represented as a set of states, environment events, transition probabilities, and state rewards [15]. In this representation, the mental model is the predicted probability that the human believes the system is in each state. Higher levels of situation awareness can be estimated by rolling out the Markov process and assuming players and opponents will perform actions that maximize their reward. Additionally, the state rewards and transition probabilities can be tuned to individual teammates to personalize the mental model to the teammate’s experience and characteristics.
The POMDP representation meets the criteria of a mental model in that it models the belief state of an agent. However, it does so by inverting the representation from an object-based belief state to a state-based belief state. In a low-dimensional environment, it is feasible to represent all possible states and their transition probabilities. However, the complex environments sought by HRI domains can overwhelm the capabilities of a POMDP. Few works have applied the POMDP representation to real-world scenarios.
Figure 1: Overview of the predicted user belief state system. The ground truth world state information is filtered by the robot’s visibility ($R^{robot})$ and used to update the robot’s belief state, $\beta^{robot}$. The robot’s belief state is then filtered by the user’s visibility to update the predicted user belief state, $\beta^{pred}$. The resulting belief state is what the robot thinks the user is aware of, i.e., a theory of mind that can inform downstream reasoning tasks.
II-C Fuzzy Mental Model Finite State Machine (FMMFSM)
Bolton et al. recently proposed FMMFSM, a model inspired by fuzzy logic concepts that uses membership functions to estimate a teammate’s belief state [2]. FMMFSM addresses the uncertainty problem of the logical predicates model, and uses an object-based scene graph. While the logical predicates model translates directly from observations to an updated mental model, FMMFSM introduces three intermediate functions:
  • $\alpha$
    maps possible environment events to the probability that each event is perceived by the teammate (addressing the observability limitation of logical predicates).
  • $\phi$
    maps perceived environment events to the probability that the event was correctly perceived.
  • $\delta$
    maps perceived environment events to the probability that the event changes each aspect of the system state.
The function implementations are arbitrary and not limited to traditional hand-crafted fuzzy membership functions.
While FMMFSM addresses key shortcomings of the logical predicates and POMDP models, it solely targets Level 1 situation awareness, or the construction of the scene graph. Predictions of contextual or future information would need to be derived from its output. As this work focuses on inference over a scene graph, instead of the scene graph itself, we chose not to utilize FMMFSM in this work. However, FMMFSM presents a compelling model that warrants evaluation in real-world human-robot teaming domains.
III Problem Formulation
Our objective is to map a set of observable environment states $O$ to a set of belief states $\{\beta^{true},\beta^{robot},\beta^{pred}\}$ over a time series. The belief states represent task-relevant world state and contextual information about the environment for an oracle ($\beta^{true}$), a robot ($\beta^{robot}$), and a human user ($\beta^{pred}$). $\beta^{true}$ is the ground truth representation of the environment, and is updated from the full environment state $O_{t}$ at each timestep $t$. $\beta^{robot}$ uses the subset of $O_{t}$ that is observable by the robot within an observable region, $R^{robot}$. Lastly, $\beta^{pred}$ is a predicted belief state of the human user from the robot’s perspective, updated using $\beta^{robot}$ and the user’s observable region $R^{user}$.
We approach the problem by defining two functions, $\Phi$ and $\mathcal{B}$, to update the belief state from observations.
$\Phi:O_{t},R\mapsto O^{\prime}_{t}$ (1)
$\Phi$ filters environment observations $O_{t}$ by a human or robot agent’s observable region $R$ to obtain the subset of observations $O^{\prime}_{t}$ that are directly observable by the agent. In a 2D domain, if $R$ is a circle of radius $r$, then $O^{\prime}_{t}$ is the set of observations from $O_{t}$ that are within $r$ from the agent. In a 3D domain, $R$ can use the agent’s pose and perception capabilities to model the agent’s visible region. When inferring belief states, such as $\beta^{pred}$, $O_{t}$ is extracted from $\beta^{robot}$ and can differ from the true world state.
$\mathcal{B}:O^{\prime}_{t},\beta_{t-1}\mapsto\beta_{t}$ (2)
$\mathcal{B}$ updates a belief state from observations. $\beta$ is the current belief state, and $O^{\prime}_{t}$ is the subset of environment observations that are used to update the belief state at a time $t$.
The core research objective of inferring a human teammate’s belief state is to identify the function $\mathcal{B}$ that maximizes the agreement between $\beta_{t}$ and the teammate’s internal belief state. In this work, we model a human user’s belief state ($\beta^{pred}_{t}$) and score the agreement between $\beta^{pred}_{t}$ and queries asked of study participants ($\beta^{human}$). In Fig. 1, $\beta^{true}$ is the ground truth of the environment, equivalent to a robot with full observability.
We evaluate the following hypotheses:
  1. H1
    A fully observable logical predicates baseline ($\mathcal{B}^{LP}$) will result in responses to situation awareness queries that moderately agree with user responses.
  2. H2
    The performance of the logical predicates baseline ($\mathcal{B}^{LP}$) at inferring user situation awareness will sharply degrade in partially observable conditions.
  3. H3
    A large language model ($\mathcal{B}^{LP+LLM}$) will match the logical predicates baseline ($\mathcal{B}^{LP}$) across fully observable and partially observable environments.
IV Methods
Our study extends the Overcooked-AI environment to predict user belief states in partial observability conditions [4]. The environment simulates a restaurant kitchen where several agents collaborate to cook and deliver soups using ingredients scattered across the kitchen’s counters. While the domain is a discretized representation of a cooking task, it presents a team-oriented structure and a fast-paced, engaging task. Additionally, the discrete nature is ideal for evaluating the logical predicates model, which the HRI community has only explored in domains that are even more simplistic [11].
In Overcooked-AI, agents in a human-robot team perform a sequence of subtasks to cook soups. Agents first move three ingredients (onions or tomatoes) from countertops to a pot. The pot then cooks for $10$ seconds, and once complete, an agent ‘plates’ the soup using a dish. Lastly, the plated soup is brought to a designated serving station, where it leaves the gameboard. The process repeats until all ingredients have been used or $90$ seconds have elapsed. Users are told to focus on completing all possible soups, but are not evaluated on their performance.
The agents share a workspace, but cannot communicate, meaning the team benefits from agents being proactive in their work and considerate of their teammate’s belief state. While related work has used reinforcement learning for the robot agent [4], our preliminary work found that a task-oriented state machine with A* planning was effective across our kitchen layouts. In this work, we are focused on predicting the user belief state ($\beta^{pred}$), not applying that prediction to planning and decision-making.
IV-A Partial Observability
We adapted the Overcooked-AI environment to simulate real-world conditions by introducing partial observability. Our visibility function, $\Phi$, limits an agent’s field of view to a region defined by an agent-specific parameter $R$. The parameter includes a type (V, O, or D) and a radius:
  • V-type is a $90^{\circ}$ conic view facing the agent’s direction, which represents a narrow field of view.
  • O-type is a $360^{\circ}$ circular view of the surrounding area, and can model full observability when the radius exceeds the environment dimensions.
  • D-type is a $180^{\circ}$ semi-circular view of the area in front of the agents, and can model half observability when the radius exceeds the environment dimensions.
Fig. 2 shows an example of each visibility type. Agents are able to see the environment objects within their visible region $R$. As we would expect real-world agents to track teammates via sound and sight, as well as have a strong knowledge of the static kitchen components, agents can always see the locations of their teammate, the kitchen floorplan, and the kitchen appliances (pot, serving station).
Figure 2: The task domain as shown to the user (green agent), with three visibility types and radii (e.g., V3 indicates V-type with radius $3$). The green shaded region is visible to the user. Each type is a unique shape representing a common field of view, and the radius indicates the extent of the visible region. For our experiment, all users had D4 visibility, and we varied the visibility of the robot (blue agent) in our post-hoc analysis.
Figure 3: Four views of the same scene. The Raw Observations view shows the direct world state $O_{t}$, the Full Observability Belief State ($\beta^{true}$) shows the ground truth beliefs, the Robot’s Partial Observability Belief State ($\beta^{robot}$) is updated from the world state within the blue region around the robot ($R^{robot}$), and the Estimated Human’s Belief State ($\beta^{user}$) is updated from the robot’s belief state filtered by the green region around the user ($R^{user}$). Grey edges indicate that two objects can be used with each other at this time. A false belief is highlighted in red: the user set a plate upon a counter, however since it occurred outside the view of the robot, the robot’s belief state and the estimated human belief state still show the plate as being held.
IV-B Object Permanence in the Scene Graph
We considered belief states as two components – a scene graph constructed from the filtered environment observations $O^{\prime}_{t}$, and an inference over the scene graph to infer contextual information. However, observations do not inherently track objects, and without object permanence the agent’s belief state cannot support objects moving in and out of the visible region. Thus, we require a function to track objects across updates to the scene graph and resolve conflicts. Object permanence is still a challenging open research problem [24].
In addition to object permanence, the cooking environment is challenging because of object transformation. When ingredients are used to make a soup, the ingredients simply no longer exist. Similarly, soups leave the gameboard when they are placed on the serving station. Our system must not only track objects through time, but also identify when objects have transformed and predict which objects were transformed. To help simplify this problem, we assume that all ingredient objects are known at the task’s initialization.
Object permanence was maintained using a structured three-pass logical approach. In the first pass, static environment objects are identified and matched to their corresponding known object. While this naive case can be tricked by swapping an item with an identical item, it is not necessary for our belief states to identify that edge condition.
The second pass matches objects to their closest known object. Similar to the naive case, the pass matches each remaining observed object to the closest remembered object of the same class. This is most effective with frequent low-ambiguity observations, which is feasible for a two-agent team. To resolve objects that were picked up by agents, the pass defaults to the closest unmatched known object.
The final pass addresses transformed objects. These are observed objects that have no fit to a known object (i.e., a newly-made soup), or known objects that have no observed match (i.e., a plated soup that was delivered). Due to the partially observable environment, it can be challenging to determine which objects transformed into other objects, however many cases can be inferred by tracking a soup’s ingredients. For example, a new “tomato, onion, and tomato soup” object implies that two known tomatoes and one known onion no longer exist on the gameboard. As each agent’s belief state is initialized with the true gameboard state, and we can assume no ingredients are added during gameplay, this pass can reasonably identify which ingredients were used to create a soup and when a soup has likely been served.
Due to the freedom users have to select ingredients and complete the task, no object permanence system can be perfect in all partially observable environments.
Fig. 3 shows a reconstruction of the gameboard using three belief states. Notably, the figure shows that the robot’s belief state $\beta^{robot}$ mistracks a plate that the user set down outside of the robot’s visible region $R$, which cascades to a misbelief in the predicted user belief state $\beta^{pred}$.
IV-C Inferring Context and Situation Awareness
The value of modeling a teammate’s belief state lies in applying the belief state to correctly infer additional contextual information and answer questions. We use the SAGAT method [8] to obtain the user’s real belief state, $\beta^{human}$, by periodically pausing the game and asking the user questions about the environment. Up to two questions are asked every $30$ seconds. We include questions about the first two levels of situation awareness – world state understanding and contextual information. All situation awareness questions were multiple choice and randomly chosen from a set question bank.
As shown in Fig. 1, the user responses are compared to the predicted belief state $\beta^{pred}$ to evaluate the performance of the belief state function $\mathcal{B}$. We compare two functions, $\mathcal{B}^{LP}$ and $\mathcal{B}^{LP+LLM}$.
For the logical predicates model, $\mathcal{B}^{LP}$, hand-crafted rules use the current scene graph to derive responses to each situation awareness question. For example, to answer the contextual information question “How many more soups can be made/delivered?”, $\mathcal{B}^{LP}$ calculates the maximum number of remaining soups using the ingredients and plates known to $\beta^{pred}$. Consequently, the logical predicates model is unable to answer questions that are not hand-crafted beforehand.
$\mathcal{B}^{LP+LLM}$ uses $\mathcal{B}^{LP}$ to construct the predicted scene graph component of $\beta^{pred}$, and an LLM (GPT4 [1]) to reason upon $\beta^{pred}$ to answer situation awareness questions. The model is prompted with a description of the game and the current scene graph representation, as well as the query “You are A1 […] Please answer the question using only one of responses below […] What is your answer?” To fairly compare $\mathcal{B}^{LP}$ and $\mathcal{B}^{LP+LLM}$, the models do not consider prior world states nor data from other participants, although exploring alternative methods of constructing the inferred scene graph is an opportunity for future work. The primary benefit of using an LLM for this task is the LLM’s ability to answer open vocabulary questions about the scene, which could supercede a logical predicates model. Additionally, an LLM can distill its reasoning into a logical format compatible with a $\mathcal{B}^{LP}$ model for efficient repeat queries.
IV-D Scoring and Evaluation
Belief states are scored by comparing the responses of the belief states to situation awareness questions. Correct responses are awarded one point, and incorrect responses zero. For questions based on spatial reasoning, we awarded partial credit for near-correct answers. For example, for a question asking “Where is the closest tomato?”, responding “Center” when the correct answer was one tile off-center scored $0.5$. The scoring function is defined as:
$Score_{u,l}=\frac{\sum\limits_{q,t\in N}(score_{q}(\beta^{A}_{t},\beta^{B}_{t})}{|N|}$ (3)
Where $q$ is a situation awareness question asked to the user $u$ on layout $l$ at time $t$, $N$ is the total number of situation awareness questions asked, and $score_{q}(\beta^{A}_{t},\beta^{B}_{t})\mapsto\mathbb{R}$ is the scoring function for $q$ using two belief states $\beta^{A}_{t}$ and $\beta^{B}_{t}$. Our scoring function was symmetric, such that $score_{q}(\beta^{A}_{t},\beta^{B}_{t})=score_{q}(\beta^{B}_{t},\beta^{A}_{t})$.
V User Study Design
We evaluated our methods with data collected through an online user study approved by Georgia Tech’s IRB. Participants were recruited from the research platform Prolific [16]. The study’s screening criteria included accessing the study using a personal computer, being within the ages of 18-89, self-identification as not part of any vulnerable group, being physically located in the United States, and having English as a first language. The study took approximately 20 minutes to complete, and participants were compensated \$$\$5$ USD.
Participants completed two practice runs and then four trials with our experiment domain. Each trial used a different kitchen layout. The game’s state was recorded at each frame ($10$hz), as well as responses to situation awareness questions.
Post-hoc, the game state histories were replayed and the logical predicates model was used to construct $\beta^{true}$, $\beta^{robot}$, and $\beta^{pred}$ across a range of visibilities for $R^{robot}$. The $\mathcal{B}^{LP}$ and $\mathcal{B}^{LP+LLM}$ models were then used to compare responses derived from the models to the actual user responses.
Figure 4: Violin plot of the score distributions of user responses to situation awareness questions ($\beta^{human}$) with respect to the ground truth oracle ($\beta^{true}$). The broad variances show the task is appropriately difficult.
VI Results & Discussion
A total of $42$ participants completed the study ($50\%$ female, ages $22-74$). We excluded $11$ participants who either failed to complete the study or failed to outperform random selection of the situation awareness responses, resulting in $31$ participants included in our evaluation.
VI-A User Performance
We first judge whether the task was appropriately difficult. Fig. 4 shows the distribution of user scores with respect to the ground truth across the four evaluated kitchen layouts. While Layout 3 appears to be easier than the other layouts, the other score distributions have similar medians, indicating consistent difficulty. Despite several perfect scores at each layout, the broad score distributions indicate that the task was neither trivial nor overwhelming for users.
Notably, few users scored perfectly (scores of $100\%$), and no users scored consistently well across all layouts. This demonstrates that users were generally unable to fully track the environment, motivating the potential benefits of predicting user belief states in this partially observable domain.
VI-B Fully Observable Robot Belief State
Hypothesis H1 evaluates the logical predicates baseline in its ideal conditions. We are interested in how accurately a robot with full observability can predict the situation awareness responses of a human user with partial observability. Unlike our evaluation of the User Performance, wherein the human user was scored against the ground truth oracle, in this case the robot does not assume that the user has any knowledge of the environment outside of what the robot knows the user has explicitly seen.
We find that the scores from the predicted user belief state ($\beta^{pred}$) with respect to the actual user responses ($\beta^{human})$ are almost identical to Fig. 4. The finding indicates that there were virtually no instances in our study where a model assuming perfect user perception could have had a false belief, i.e., environment changes that would have impacted the situation awareness responses were quickly observed by users, although seemingly not always recalled.
Additionally, Fig. 2 shows that the full observability model’s performance averaged between $55\%$ and $75\%$. This suggests that users had nuanced perception, comprehension, or memory that was not captured by $\mathcal{B}^{LP}$. Therefore, opportunities exist for alternative models to utilize human factors that can contribute to a user’s belief state.
Thus, we affirm our hypothesis H1.
Figure 5: Line plot of the mean performance of $\beta^{pred}$ using $\mathcal{B}^{LP}$ with respect to the user responses, across various visibility parameters $R^{robot}$. While performance declines between the three visibility types, there is not a meaningful decline between visibility conditions.
VI-C Partial Robot Observability
Hypothesis H2 addresses how $\mathcal{B}^{LP}$ affects the performance of $\beta^{robot}$ across a range of visibilities. For a belief state function $\mathcal{B}$ to be feasible in human-robot teaming domains, it is valuable to see how the function’s performance declines with decreasing visibility.
Fig. 5 shows the performance of $\mathcal{B}^{LP}$ across $14$ visibility conditions. Counter to our hypothesis H2, the performance of $\mathcal{B}^{LP}$ only slightly decreased between high visibility and low visibility conditions. We believe this due to our logical predicates implementation accurately resolving conflicting beliefs during the task. The findings support using $\mathcal{B}^{LP}$ as a baseline for more complex partially observable domains.
Thus, we refute our hypothesis H2.
Figure 6: Chart of the mean performances of $\beta^{pred}$ constructed using the $\mathcal{B}^{LP}$ and $\mathcal{B}^{LP+LLM}$ models over two observability conditions. Error bars show the variance. $\mathcal{B}^{LP+LLM}$ has comparable means and variances to $\mathcal{B}^{LP}$ when scored against user responses, indicating that $\mathcal{B}^{LP+LLM}$ is able to infer an inference ruleset similar to the hand-crafted developed for $\mathcal{B}^{LP}$. Higher scores indicate superior performance.
VI-D Large Language Model
Hypothesis H3 explores whether an LLM can predict and reason about the user’s belief state as a zero-shot system. The open vocabulary reasoning capability of LLMs has no equivalent for logical predicates, and would be valuable for broadly applying predicted belief states. As a first step, we are interested in whether an LLM can match the performance of the logical predicates model ($\mathcal{B}^{LP}$).
Fig. 6 shows the performance of $\mathcal{B}^{LP}$ and $\mathcal{B}^{LP+LLM}$ at predicting user responses to situation awareness questions. The two models were presented with the same scene graph and prompted with which answer the user would choose, and we found no meaningful difference in mean nor variance between the two models. The findings support the use of LLMs for predicting user belief states, and as a baseline upon which to compare future models.
Thus, we affirm our hypothesis H3.
VII Conclusion
For decades, the logical predicates model has been a de-facto representation of world states, and has recently been extended to represent belief state prediction for human-robot teaming. The logical structure makes the model compatible with state machines, linear temporal logic planners, and case-based reasoning, which is attractive for real-time systems seeking transparent knowledge representations.
However, the logical predicates model does not consider the probabilistic nature of human cognition – people do not see the world in absolute binary truths, but as probability and managed uncertainty. As such, the logical predicates model is challenged to represent individual human differences in their interpretations of the environment. Moreover, very few works have operationalized the model to human-robot teaming domains, and the literature typically assumes that robot agents have full observability over their environments.
Unable to consider uncertainty, the logical predicates model is limited to low-dimension state spaces and high-certainty predicates. As human-robot teaming domains typically operate in complex environments, misinterpretations can have consequences on downstream planning.
In this work, we explored the strengths and limitations of using a logical predicates belief state model in human-robot teaming. We defined the problem of predicting a co-located human teammate’s belief state, measured by the model’s responses to situation awareness questions. Through a user study we collected a dataset of responses at various points throughout a collaborative cooking game, which we utilized to test the accuracy of a logical predicates model and an LLM at predicting the user’s responses under several degrees of full and partial observability.
Key takeaways from this work include:
  1. 1.
    The logical predicates model resulted in a scene graph where neither inference model scored above $50-75\%$, motivating the integration of human factors techniques to construct predicted scene graphs for teammates.
  2. 2.
    The performance of the logical predicates model only slightly degraded between fully observable and partially observable conditions, supporting the model’s use in more realistic partially observable domains if a careful implementation is feasible.
  3. 3.
    The performance of the logical predicates model was matched by a large language model in a zero-shot manner, suggesting that LLMs can be leveraged to help construct logical predicate models, or even supercede them in more complex environments.